# Stochastic Vector Approximate Message Passing with applications to phase retrieval

### v1

@@ 1.1 Background on Phase Retrieval
It is of practical interest how many observations are needed to recover the signal $\boldsymbol{x}$. In recent years, the information-theoretic threshold of sampling ratio has been intensively studied for phase retrieval with large random sensing matrices [2]. From an algorithmic viewpoint, spectral method  [3, 4, 5] is known to be an efficient algorithm for obtaining an estimate $\hat{\boldsymbol{x}}$ correlated with $\boldsymbol{x}$, which is often employed to initialize iterative solvers, including the Wirtinger flow [6], and message passing algorithms [7].
@@
Regarding prior knowledge of the object image, support constraint and sparsity of $\boldsymbol{x}$ are frequently exploited in phase retrieval. Bayesian inference provides a framework for incorporating such prior knowledge into a statistical model. The Bayes’ theorem for the model (1) is written as
@@
wherein the prior $p_{\text{in}}(\boldsymbol{x})$ expresses the prior knowledge on $\boldsymbol{x}$, and $\delta(\cdot)$ is the Dirac delta function. To compute the posterior density $p(\boldsymbol{x},\boldsymbol{z}|\boldsymbol{y})$ in a reasonable computational time, approximation methods such as the mean-field approach and approximate message passing have been successfully applied to phase retrieval. [2, 11, 12]
@@ 1.2 Illustration of the key idea
To incorporate multiple measurements into a Bayesian framework, we extend Vector Approximate Message Passing (VAMP)  [15, 16], an efficient inference algorithm for generalized linear models. Our extension, coined Stochastic VAMP, introduces sequential access to data subsets, significantly speeding up convergence. This method is analogous to Stochastic Gradient Descent in non-convex optimization, where gradients are computed from data subsets, hence the name Stochastic VAMP.
@@ 2.1 Derivation of VAMP
$\boldsymbol{x}_{1}=\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})\ ,\ \eta_{1}^{-1}=\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1},\gamma_{1})$ (8)
@@
Likewise, the message $\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p},\tau_{1}^{-1}I_{M})$ is input to $\boldsymbol{z}_{1}$ node. $\hat{\boldsymbol{z}}_{1},\lambda_{1},\boldsymbol{p}_{2}$, and $\tau_{2}$ are give by
@@
$\boldsymbol{z}_{1}=\boldsymbol{g}_{z}(\boldsymbol{p}_{1},\tau_{1})\ ,\ \lambda_{1}^{-1}=\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}_{1},\tau_{1})$ (12)
@@
with $b(\boldsymbol{z}_{1})\propto\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p},\tau_{1}^{-1}I_{M})p_{\text{out}}(\boldsymbol{y}||\boldsymbol{z}_{1}|)$.
@@ 2.2 Stochastic VAMP
2.3 Algorithmic detail
We note some implementation details of Stochastic VAMP. Typically, damping is used in the denoising step to help VAMP converge, with the update rule $\hat{\boldsymbol{x}}_{1k}=\rho\boldsymbol{g}_{x}(\boldsymbol{r}_{1k},\gamma_{1k})+(1-\rho)\hat{\boldsymbol{x}}_{1(k-1)}$ and so on, where $10<\rho<$ 1. For unstructured matrix such as Haar distributed $A$ in Sec.3.2, Stochastic VAMP does not need damping. For the sensing matrices of coded diffraction pattern, a small amount of damping is helpful. In Sec.3.3, we used damping parameter $\rho=0.95$.
Secondly, when each matrix $A^{(l)}$ have $(A^{(l)})^{H}A^{(l)}$ diagonal, the matrix inversion in line 5 adds only minor complexity to the algorithm. The dominant computational cost comes from the matrix-vector multiplications in lines 6 and 7. In the case of coded diffraction patterns, the main computational cost is $2L$ FFTs per iteration. For general sensing matrices with non-diagonal $A^{H}A$, costly matrix-inversion can still be avoided by using the graphical model introduced for D-VAMP [18] and pre-computing the singular value decompositions of $A^{(1)},...,A^{(L)}$.
Lastly, in our experiments, we assumed that the SNR and prior are known, but they can be learned from the dataset using methods from  [19]. Although noise learning and multiple restarts can help phase retrieval algorithms avoid bad local minima [11, 12], numerical results indicate that Stochastic VAMP remains stable even without these enhancements.
@@ 2.3 Algorithmic detail
1:Sensing matrices $A^{(1)},...,A^{(L)}$, observations $\boldsymbol{y}^{(1)},...,\boldsymbol{y}^{(L)}$, number of iterations $K_{it}$
2:Initialize $\boldsymbol{r}_{20},\gamma_{20}>0$ and $\boldsymbol{p}_{20},\tau_{20}>0\ \ (l=1,...,L)$
3:for $k=0,1,...,K_{it}$ do
@@
6:         $Q=\left(\gamma_{2k}I_{N}+\sum_{l=1}^{L}\tau^{(l)}_{2k}(A^{(l)})^{H}A^{(l)}\right)^{-1}$
7:         $\hat{\boldsymbol{x}}_{2k}=Q\left(\gamma_{2k}\boldsymbol{r}_{2k}+\sum_{l=1}^{L}\tau^{(l)}_{2k}(A^{(l)})^{H}\boldsymbol{p}^{(l)}_{2k}\right)$
8:         $\eta_{2k}^{-1}=\frac{1}{N}\text{Tr}(Q)$
9:         $\hat{\boldsymbol{z}}_{2k}=A^{(l)}\hat{\boldsymbol{x}}_{2k}$
10:         $(\lambda^{(l)}_{2k})^{-1}=\frac{1}{M}\text{Tr}\left(A^{(l)}Q(A^{(l)})^{H}\right)$
@@
12:         $\gamma_{1k}=\eta_{2k}-\gamma_{2k}$
13:         $\boldsymbol{r}_{1k}=(\eta_{2k}\hat{\boldsymbol{x}}_{2k}-\gamma_{2k}\boldsymbol{r}_{2k})/\gamma_{1k}$
14:         $\tau_{1k}^{(l)}=\lambda_{2k}^{(l)}-\tau_{2k}^{(l)}$
15:         $\boldsymbol{p}^{(l)}_{1k}=(\lambda_{2k}^{(l)}\hat{\boldsymbol{z}}_{2k}-\tau_{2k}^{(l)}\boldsymbol{p}^{(l)}_{2k})/\tau_{1k}^{(l)}$
@@
17:         $\hat{\boldsymbol{x}}_{1k}=\boldsymbol{g}_{x}(\boldsymbol{r}_{1k},\gamma_{1k})$
18:         $\eta_{1k}^{-1}=\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1k},\gamma_{1k})$
19:         $\hat{\boldsymbol{z}}^{(l)}_{1k}=\boldsymbol{g}_{z}(\boldsymbol{p}^{(l)}_{1k},\tau_{1k})$
20:         $\lambda_{1k}^{-1}=\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}^{(l)}_{1k},\tau_{1k})$
@@
22:         $\gamma_{2(k+1)}=\eta_{1k}-\gamma_{1k}$
23:         $\boldsymbol{r}_{2(k+1)}=(\eta_{1k}\hat{\boldsymbol{x}}_{1k}-\gamma_{1k}\boldsymbol{r}_{1k})/\gamma_{2(k+1)}$
24:         $\tau_{2(k+1)}^{(l)}=\lambda_{1k}^{(l)}-\tau_{1k}^{(l)}$
25:         $\boldsymbol{p}^{(l)}_{2(k+1)}=(\lambda_{1k}^{(l)}\hat{\boldsymbol{z}}_{1k}-\tau_{1k}^{(l)}\boldsymbol{p}^{(l)}_{1k})/\tau_{2(k+1)}^{(l)}$
@@
28:Return $\hat{\boldsymbol{x}}_{1k}$
@@ 3.1 Common setup
where $\boldsymbol{w}$ is the additive white Gaussian noise (AWGN) given as $\boldsymbol{w}\sim\mathcal{N}(\cdot|\boldsymbol{0},\gamma_{w}^{-1}I_{M})$, and ||$\cdot$|| ⋅ | is the element-wise absolute value. The corresponding likelihood function $p_{\text{out}}(\boldsymbol{y}|\boldsymbol{z})$ is refered to as the Rician distribution, whose denoising functions are derived in  [12].
We assume a Gaussian prior $p_{\text{in}}(\boldsymbol{x})=\mathcal{N}(\boldsymbol{x}|\boldsymbol{0},I_{N})$, but structured priors such as sparse priors can also be used. The total number of observed values is denoted as $M=L\bar{M}$, where $L$ is the number of sensing matrices $A^{(1)},...,A^{(L)}$ and $A^{(l)}\in\mathbb{C}^{\bar{M}\times N}$. The sampling ratio is $\alpha\triangleq\frac{M}{N}$.
@@ 3.2 Haar distributed sensing matrix
We consider sensing matrices $A^{(l)}$ with $(A^{(l)})^{H}A^{(l)}=I_{\bar{M}}$ drawn from Haar distribution. In the case of noiseless phase retrieval, the threshold value of $\alpha$ above which VAMP recovers the true signal $\boldsymbol{x}$ is investigated via asymptotic analysis [2], in which the threshold value for Haar-distributed $A$ is estimated as $\alpha_{\text{FR, Alg}}\simeq 2.265$.
In our experiment, we used $N=512$, $\alpha=2.4$, and $L=2$, with a noise level of 30 dB. We compared Stochastic VAMP with vanilla VAMP, which uses concatenated sensing matrices and observations as in Eq.(4). While VAMP required damping ($\rho=0.7$) for convergence, no damping was needed for Stochastic VAMP. The true signal $\boldsymbol{x}$ was drawn from the prior, and both algorithms were randomly initialized. The complexity of both algorithms is similar, with $O((L-1)N)$ difference per-iteration. The NMSE $\frac{||\boldsymbol{x}-\hat{\boldsymbol{x}}_{1k}||^{2}}{||\boldsymbol{x}||^{2}}$ is shown against iteration number $k$ in Fig. 4. We observed that the stochasticity significantly accelerates the convergence of the phase retrieval algorithm.

### v3

@@ 1.1 Background on Phase Retrieval
It is of practical interest how many observations are needed to recover the signal $\boldsymbol{x}$. In recent years, the information-theoretic threshold of sampling ratio has been intensively studied for phase retrieval with large random sensing matrices [2]. From an algorithmic viewpoint, spectral method  [3, 4] is known to be an efficient algorithm for obtaining an estimate $\hat{\boldsymbol{x}}$ correlated with $\boldsymbol{x}$, which is often employed to initialize iterative solvers, including the Wirtinger flow [5], and message passing algorithms [6].
@@
Regarding the latter approach, Bayesian inference provides a framework for incorporating prior knowledge, such as sparsity and support constraint, into a statistical model. The Bayes’ theorem for the model (1) is written as
@@
wherein the prior $p_{\text{in}}(\boldsymbol{x})$ expresses the prior knowledge on $\boldsymbol{x}$, and $\delta(\cdot)$ is the Dirac delta function. To compute the posterior density $p(\boldsymbol{x},\boldsymbol{z}|\boldsymbol{y})$ in a reasonable computational time, approximations such as the mean-field approach and approximate message passing have been applied to phase retrieval [2, 10, 11]. These techniques can also be combined with deep neural network trained with natural images [12].
@@ 1.2 Illustration of the key idea
To incorporate multiple measurements into a Bayesian framework, we extend Vector Approximate Message Passing (VAMP)  [15, 16], an efficient inference algorithm for generalized linear models. Our extension, coined Stochastic VAMP, allows ”mini-batch learning” commonly employed in ptychographic reconstruction[14] and speeds up convergence. This method is analogous to Stochastic Gradient Descent in non-convex optimization, where gradients are computed from data subsets, hence the name Stochastic VAMP.
@@ 2.1 Derivation of VAMP
$\hat{\boldsymbol{x}}_{1}=\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})\ ,\ \eta_{1}^{-1}=\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1},\gamma_{1})$ (8)
@@
Likewise, the message $\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p}_{1},\tau_{1}^{-1}I_{M})$ is input to $\boldsymbol{z}_{1}$ node. $\hat{\boldsymbol{z}}_{1},\lambda_{1},\boldsymbol{p}_{2}$, and $\tau_{2}$ are give by
@@
$\hat{\boldsymbol{z}}_{1}=\boldsymbol{g}_{z}(\boldsymbol{p}_{1},\tau_{1})\ ,\ \lambda_{1}^{-1}=\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}_{1},\tau_{1})$ (12)
@@
with $b(\boldsymbol{z}_{1})\propto\mathcal{N}(\boldsymbol{z}_{1}|\boldsymbol{p}_{1},\tau_{1}^{-1}I_{M})p_{\text{out}}(\boldsymbol{y}||\boldsymbol{z}_{1}|)$.
@@ 2.2 Stochastic VAMP
1:Sensing matrices $A^{(1)},...,A^{(L)}$, observations $\boldsymbol{y}^{(1)},...,\boldsymbol{y}^{(L)}$, number of iterations $K_{\text{it}}$
2:Initialize $\boldsymbol{r}_{2},\gamma_{2}>0$ and $\boldsymbol{p}_{2},\tau_{2}>0\ \ (l=1,...,L)$
3:for $k=0,1,...,K_{\text{it}}$ do
@@
6:         $Q=\left(\gamma_{2}I_{N}+\sum_{l=1}^{L}\tau^{(l)}_{2}(A^{(l)})^{H}A^{(l)}\right)^{-1}$
7:         $\hat{\boldsymbol{x}}_{2}=Q\left(\gamma_{2}\boldsymbol{r}_{2}+\sum_{l=1}^{L}\tau^{(l)}_{2}(A^{(l)})^{H}\boldsymbol{p}^{(l)}_{2}\right)$
8:         $\eta_{2}^{-1}=\frac{1}{N}\text{Tr}(Q)$
9:         $\hat{\boldsymbol{z}}_{2}=A^{(l)}\hat{\boldsymbol{x}}_{2}$
10:         $(\lambda^{(l)}_{2})^{-1}=\frac{1}{M}\text{Tr}\left(A^{(l)}Q(A^{(l)})^{H}\right)$
@@
12:         $\gamma_{1}=\eta_{2}-\gamma_{2}$
13:         $\boldsymbol{r}_{1}=(\eta_{2}\hat{\boldsymbol{x}}_{2}-\gamma_{2}\boldsymbol{r}_{2})/\gamma_{1}$
14:         $\tau_{1}^{(l)}=\lambda_{2}^{(l)}-\tau_{2}^{(l)}$
15:         $\boldsymbol{p}^{(l)}_{1}=(\lambda_{2}^{(l)}\hat{\boldsymbol{z}}_{2}-\tau_{2}^{(l)}\boldsymbol{p}^{(l)}_{2})/\tau_{1}^{(l)}$
@@
17:         $\hat{\boldsymbol{x}}_{1}=\boldsymbol{g}_{x}(\boldsymbol{r}_{1},\gamma_{1})$
18:         $\eta_{1}^{-1}=\boldsymbol{g}^{\prime}_{x}(\boldsymbol{r}_{1},\gamma_{1})$
19:         $\hat{\boldsymbol{z}}^{(l)}_{1}=\boldsymbol{g}_{z}(\boldsymbol{p}^{(l)}_{1},\tau_{1})$
20:         $\lambda_{1}^{-1}=\boldsymbol{g}^{\prime}_{z}(\boldsymbol{p}^{(l)}_{1},\tau_{1})$
@@
22:         $\gamma_{2}=\eta_{1}-\gamma_{1}$
23:         $\boldsymbol{r}_{2}=(\eta_{1}\hat{\boldsymbol{x}}_{1}-\gamma_{1}\boldsymbol{r}_{1})/\gamma_{2}$
24:         $\tau_{2}^{(l)}=\lambda_{1}^{(l)}-\tau_{1}^{(l)}$
25:         $\boldsymbol{p}^{(l)}_{2}=(\lambda_{1}^{(l)}\hat{\boldsymbol{z}}_{1}-\tau_{1}^{(l)}\boldsymbol{p}^{(l)}_{1})/\tau_{2}^{(l)}$
@@
28:Return $\hat{\boldsymbol{x}}_{1}$
2.3 Algorithmic detail
It is standard to stabilize the convergence of VAMP by employing the ”damping” method, where $\boldsymbol{r}_{1}$ in Eq. (17) is treated as ”raw” value $\boldsymbol{r}_{\text{raw}}$, and $\boldsymbol{r}_{1}$ is updated as $\boldsymbol{r}_{1}=\rho\boldsymbol{r}_{\text{raw}}+(1-\rho)\boldsymbol{r}_{\text{old}}$, in which $10<\rho<$ 1 and $\boldsymbol{r}_{\text{old}}$ is the previous value of $\boldsymbol{r}_{1}$. Similar technique is applied to the update of $\gamma,\boldsymbol{p},$ and $\tau$. Notably, Stochastic VAMP requires less damping than standard VAMP, contributing to a faster convergence.
When each matrix $A^{(l)}$ have $(A^{(l)})^{H}A^{(l)}$ diagonal, the matrix inversion in line 5 adds only minor complexity to the algorithm. The dominant computational cost comes from the matrix-vector multiplications in lines 6 and 8. In the case of coded diffraction patterns, the cost is $2L$ FFTs per iteration. For general sensing matrices, costly matrix-inversion can still be avoided by using the graphical model introduced for D-VAMP [18] and pre-computing the singular value decompositions of $A^{(1)},...,A^{(L)}$.
In our experiments, we assumed that the SNR and prior are known, but they can be learned from the dataset using methods from  [19]. Although noise learning and multiple restarts can help phase retrieval algorithms avoid bad local minima [10, 11], numerical results indicate that Stochastic VAMP remains stable even without these enhancements. Python codes for the numerical experiments in this paper is available at https://github.com/sacbow/StochasticVAMP.
@@ 3.1 Common setup
where $\boldsymbol{w}$ is the Gaussian noise given as $\boldsymbol{w}\sim\mathcal{N}(\cdot|\boldsymbol{0},\gamma_{w}^{-1}I_{M})$, and ||$\cdot$|| ⋅ | is the element-wise absolute value. The corresponding likelihood function $p_{\text{out}}(\boldsymbol{y}||\boldsymbol{z}|)$ is refered to as the Rician distribution, whose denoising functions are derived in  [11]. We assume a Gaussian prior $p_{\text{in}}(\boldsymbol{x})=\mathcal{N}(\boldsymbol{x}|\boldsymbol{0},I_{N})$, but structured priors can also be used. The total number of observed values is $M=L\bar{M}$, where $L$ is the number of sensing matrices $A^{(1)},...,A^{(L)}$ and $A^{(l)}\in\mathbb{C}^{\bar{M}\times N}$. The sampling ratio is $\alpha\triangleq\frac{M}{N}$. Since the posterior density is invariant w.r.t. global phase shift, the deviation between the true signal $\boldsymbol{x}$ and VAMP estimate $\hat{\boldsymbol{x}}$ is measured by
$\text{NMSE}\triangleq\underset{\theta\in[0,2\pi)}{\operatorname{min}}\frac{||\boldsymbol{x}-e^{j\theta}\hat{\boldsymbol{x}}||^{2}}{||\boldsymbol{x}||^{2}}$ (20)
@@ 3.2 Haar distributed sensing matrix
We consider sensing matrices $A^{(l)}$ with $(A^{(l)})^{H}A^{(l)}=I_{N}$ drawn from Haar distribution. In the case of noiseless phase retrieval, the threshold value of $\alpha$ above which VAMP recovers the true signal $\boldsymbol{x}$ is investigated via asymptotic analysis [2], in which the threshold value for Haar-distributed $A$ is estimated as $\alpha_{\text{FR, Alg}}\simeq 2.265$. In our experiment, we used $N=512$, $\alpha=2.4$, and $L=2$, with SNR of 30 dB. We compared Stochastic VAMP with vanilla VAMP, which uses concatenated sensing matrices and observations (Eq.(4)). While VAMP required damping of $\rho=0.9$ for convergence, Stochastic VAMP was stable with less amount of damping ($\rho=0.97$). The true signal $\boldsymbol{x}$ was drawn from the prior, and both algorithms were randomly initialized. The complexity of both algorithms is similar, with $O((L-1)N)$ difference per-iteration. The NMSE is shown against iteration number $k$ in Fig. 4. We observed that the stochasticity significantly accelerates the convergence of the phase retrieval algorithm.